OpenStudy (anonymous):

A rainstorm in Portland, Oregon, wiped out the electricity in 5% of the households in the city. Suppose that a random sample of 50 Portland households is taken after the rainstorm. A.) Estimate the number of households in the sample that lost electricity by giving the mean of the relevant distribution (that is, the expectation of the relevant random variable). Do not round your response. B.) Quantify the uncertainty of your estimate by giving the standard deviation of the distribution. Round your response to at least three decimal places.

OpenStudy (anonymous):

I *think* i understand #A, and I got 2.5 for that answer.. but #B is where I'm having a lot of trouble.

OpenStudy (perl):

this is a binomial distribution A) 50 * .05 B) s.d = sqrt( 50 * .05 * .95)

OpenStudy (perl):

it is binomial because youre respondants in your survey either 'lost electricity' or 'did not lose electricity'

OpenStudy (anonymous):

\[\sqrt{50*.05*.95}\]? I tried that, I did 50*.05*.95 and got .2375 but there is no square root of .2375 i guess?

OpenStudy (anonymous):

having u tried pressing the calculator ? sqrt(2.375) = 1.541 (3.d.p)

OpenStudy (anonymous):

i did haha, sorry. it came up with "invalid". Thank you.

OpenStudy (anonymous):

I just tried again and it's giving me 15.3948.. is the calculator I'm using on the fritz?

OpenStudy (anonymous):

are you using TI series ?

OpenStudy (anonymous):

HEHE I live in portland OR

OpenStudy (anonymous):

I was using www.mathway.com

OpenStudy (anonymous):

yea i got 15.3948. from that website too :/

OpenStudy (anonymous):

uhm.. guess you have to use other calculator, that website is not very reliable ><

OpenStudy (anonymous):

...I'm going to stop using that website, then. Lol. Thanks so much for your help.

OpenStudy (perl):

just type into google : sqrt( 50 * 0.5 * .95) they have a calculator

OpenStudy (perl):

just type into google search bar: sqrt( 50 * 0.5 * .95)

OpenStudy (anonymous):

Whoa guys. you are doing something wrong!

OpenStudy (anonymous):

khubbz:'' \sqrt(50∗.05∗.95) ? I tried that, I did 50*.05*.95 and got .2375 but there is no square root of .2375 i guess?'' that's the reason you should put zeros in front of decimal points! The next post is: HatcrewS having u tried pressing the calculator ? sqrt(2.375) = 1.541 (3.d.p) The decimal point has changed. HatcrewS's calculator rounds to three decimal points. Fix it, and you're pretty much done.

OpenStudy (anonymous):

The answers I have are 2.5 and 1.541.. is that wrong?

OpenStudy (anonymous):

I think that it's right.